# Synergistic Integration of Coordinate Network and Tensorial Feature for Improving Neural Radiance Fields from Sparse Inputs

### v1

@@ Abstract
The multi-plane representation has been highlighted for its fast training and inference across static and dynamic neural radiance fields. This approach constructs relevant features via projection onto learnable grids and interpolating adjacent vertices. However, it has limitations in capturing low-frequency details and tends to overuse parameters for low-frequency features due to its bias toward fine details, despite its multi-resolution concept. This phenomenon leads to instability and inefficiency when training poses are sparse. In this work, we propose a method that synergistically integrates multi-plane representation with a coordinate-based network known for strong bias toward low-frequency signals. The coordinate-based network is responsible for capturing low-frequency details, while the multi-plane representation focuses on capturing fine-grained details. We demonstrate that using residual connections between them seamlessly preserves their own inherent properties. Additionally, the proposed progressive training scheme accelerates the disentanglement of these two features. We empirically show that the proposed method achieves comparable results to explicit encoding with fewer parameters, and particularly, it outperforms others for the static and dynamic NeRFs under sparse inputs.
@@ 1 Introduction
Approaches explicitly parameterizing spatial attributes through voxel-grid, hashgrid, and multi-plane have been introduced (Müller et al., 2022; Chen et al., 2022; Chan et al., 2022). These methods dramatically reduce training time and produce cleaner and more realistic images, meanwhile demanding excessive memory consumption (Lee et al., 2023). The recent works found that those representations struggle with low-frequency detail and overfit to high-frequency signals, especially when applying for dynamic scenes (Fridovich-Keil et al., 2023; Cao & Johnson, 2023; Peng et al., 2023). While those have marginally had success in the reconstruction of NeRF with the assistance of denoising penalties like total variation (Sun et al., 2023; Fridovich-Keil et al., 2023), they still lack adequate representation of low-frequency spectral features like object shapes and dynamic motion, as shown in 1(a).
To alleviate this issue, we introduce a simple yet powerful approach to fundamentally improve the performance of static and dynamic NeRFs from sparse inputs. In this framework, coordinate-based features handle low-frequency context, while multiple-plane features capture fine-grained details aligning with the spectral bias mentioned earlier. This approach yields three main benefits. First, aligning with the distinct spectral biases of heterogeneous features results in a model less sensitive to hyperparameters and performance variations related to the scene. This is achieved by avoiding both underfitting and overfitting. Second, it allows for stable training through gradual changes in their spectral biases, as discussed in the works by Lin et al. (2021); Yang et al. (2023). Lastly, it facilitates efficient parameter allocation by replacing the need for a low-resolution grid with coordinate-based features.
@@
  • We reduce the number of parameters by skipping the allocation of a spatial low-resolution grid and instead replacing it with coordinate-based features. Despite this reduction, the proposed method still shows impressive improvement compared to baselines.
@@ 2 Related Work
Early efforts incorporated pre-trained networks trained on large datasets to compensate for the lack of training data (Jain et al., 2021; Yu et al., 2021; Wang et al., 2021). Another alternative approach incorporated additional information, such as depth or color constraints, to ensure the preservation of 3D coherence (Deng et al., 2022; Yuan et al., 2022; Roessle et al., 2022; Truong et al., 2023). Without the assistance of off-the-shelf models and additional, this line of works devised new regularization to train NeRFs with fewer than ten views. Reg-NeRF incorporates patch-wise geometry and appearance regularization (Niemeyer et al., 2022). This paper verified that their regularization performs well on forward-facing examples like DTU and LLFF datasets. They did not validate object-facing scenes because this assumption demands a high correlation between adjacent views. Recently, progressively manipulating the spectrum of positioning encoding from low to high frequency proved effective in mitigating over-fitting without relying on additional information (Yang et al., 2023; Song et al., 2023). Compared to explicit representations, those still suffer from unsatisfactory visual quality, characterized by blurry boundaries. Recent studies using total variation regularization on explicit representations get rid of artifacts and construct smoother surfaces (Cao & Johnson, 2023; Fridovich-Keil et al., 2023; Sun et al., 2023). However, our findings indicate that this regularization can introduce artificial details that seem real but are not in the data. This can also result in the model failing to converge in certain scenes. We present this problem in the experimental results, both qualitatively and quantitatively.
@@
A few works attempted to use explicit parameterization with the sinusoidal encoding of coordinates, but their direction differs from our method since they mainly focus on enriching available features or fewer parameterization, as well as they did not demonstrate the role of tri-planes and coordinate features (Wang et al., 2023b; Peng et al., 2023; Lee et al., 2023). In this paper, our new approach proposes incorporating two distinct features: coordinate-based and multiple-plane features. We emphasize that the disentanglement of these two heterogeneous features is crucial for reliably constructing NeRFs in sparse inputs. The proposed method performs well even with higher-dimensional targets like dynamic NeRFs and extremely limited sparse inputs.
Figure 2: Conceptual illustration of the proposed method utilizing global contexts by coordinate networks and fine-grained details by multi-plane encoding. This method effectively displays two heterogeneous features. The number 1 indicates the use of coordinate network alone, while the symbol 1$+$2 means the use of both coordinate network and multi-plane representation.
@@ 3.1 Architecture
where, $\{W_{l},b_{l}\}_{l=1}^{L}$ are the weights and biases of the $l$-layer MLP. The subsequent process is defined in MLP $\phi_{l}(\cdot)$, where $l>2$, contains one pair of weights and biases. The output layers use different activations, such as the softplus function for density and the sigmoid function for color. The proposed residual connection allows the network to robustly maintain low-frequency spectral bias from coordinate networks without interference from multi-plane features. Our empirical findings demonstrate that this operation promotes the disentanglement of two features, aligning with a spanning diverse spectrum. A detailed analysis of this residual connection is provided in subsection 4.4 along with the performance gain by this architecture.
@@ 4 Experiments
To answer those questions, we conduct vast experiments over scenarios of two sparse input cases: a few-shot static case and a 4-dimensional dynamic case. To provide a clearer insight into the role of low-frequency representation, we include 2D image regression as an illustrative example. We also include ablation studies to substantiate the rationale behind the architectural choices in our proposed model.
We choose the datasets as in-ward-facing object poses, as they are more likely to be occluded by the objects from various viewing locations than forward-facing poses. For performance evaluation, we employ the PSNR metric to gauge image reconstruction quality. In addition, SSIM and LPIPS scores are reported to assess the perceptual quality of the rendered images. Further experimental details are described in Appendix D.
@@ 4.1 Motivation Example: Image Regression
We start by demonstrating the diversity in spectrum information the proposed method possesses in Image Regression tasks. We compare our method with FFN (Tancik et al., 2020) as a sinusoidal embedding and iNGP (Müller et al., 2022), K-Planes (Fridovich-Keil et al., 2023) as explicit parameterization. Additionally, we include CAM (Lee et al., 2023), a similar approach that combines coordinate-based networks and grid-based representation for efficient parameterization. Figure 4 shows how baselines handle low and high freqeuncy or resolution features by presenting the average magnitude spectrum by Fourier feature transform. The sinusoidal method faces an underfitting issue as it overly focuses on low frequency details, while explicit representations like iNGP and K-Planes tend to interpret low resolution features focusing on high frequency detail even when only low-resolution features are activated. Surprisingly, CAM, despite incoporating various spectral sinusodial embeddings, also struggles to capture low frequency details. This implies explicit representation such as grid or plane cannot effectively handle low frequency details without careful designs.
@@ 5 Conclusion
In this paper, we introduce refined tensorial radiance fields that seamlessly incorporate coordinate networks. The coordinate network enables the capture of global context, such as object shapes in the static NeRF and dynamic motions in the dynamic NeRF dataset. This property allows multi-plane encoding to focus on describing the finest details.
Impact Statements
@@ C.1 Hyper-parameters on the Static NeRF
The proposed model incorporates multi-plane encoding and MLPs with skip connections. For the multi-plane encoding, we utilize 48-dimensional channels. The resolution of plane features is upscaled to 8,000,000 ($200^{3}$) by the end of training. The hyper-parameters, such as the weight for Laplacian smoothing ($\lambda_{1}$), the curriculum learning schedule, and the initial feature resolution, vary across scenes as part of our hyperparameter tuning to achieve optimal results. Detailed information on the hyperparameters for multi-plane encoding can be found in Table C.1. For the decoder, we employ standard fully connected layers with ReLU activations, each containing 256 channels. The encoder consists of four fully connected ReLU layers, with a skip connection introduced after the second layer, which concatenates the fused input features. Occupancy is directly calculated from the obtained features with softplus function applied to the first channel. The RGB decoder, following this, consists of two layers. The color values are obtained from the features processed by the RGB decoder through sigmoid activation.
@@ C.2 Hyper-parameters on the Dynamic NeRF
The configuration for the dynamic Neural Radiance Fields (NeRF) case adheres to the same settings as the static case. We utilize plane features with 48 channels. The initial voxel resolution is set at 4,096 ($16^{3}$) and is subsequently upscaled to 8,000,000 ($200^{3}$). For a more detailed description, please refer to Table C.2. The structure of the decoder, initial learning rate, and optimizer configuration remain identical to those used in the static NeRF. Any configurations not specified here follow directly from the HexPlane method as described in (Cao & Johnson, 2023).

### v3

@@ Abstract
The multi-plane representation has been highlighted for its fast training and inference across static and dynamic neural radiance fields. This approach constructs relevant features via projection onto learnable grids and interpolating adjacent vertices. However, it has limitations in capturing low-frequency details and tends to overuse parameters for low-frequency features due to its bias toward fine details, despite its multi-resolution concept. This phenomenon leads to instability and inefficiency when training poses are sparse. In this work, we propose a method that synergistically integrates multi-plane representation with a coordinate-based MLP network known for strong bias toward low-frequency signals. The coordinate-based network is responsible for capturing low-frequency details, while the multi-plane representation focuses on capturing fine-grained details. We demonstrate that using residual connections between them seamlessly preserves their own inherent properties. Additionally, the proposed progressive training scheme accelerates the disentanglement of these two features. We demonstrate empirically that our proposed method not only outperforms baseline models for both static and dynamic NeRFs with sparse inputs, but also achieves comparable results with fewer parameters.
@@ 1 Introduction
Approaches explicitly parameterizing spatial attributes through voxel-grid, hashgrid, and multi-plane have been introduced (Müller et al., 2022; Chen et al., 2022; Chan et al., 2022). These methods dramatically reduce training time and produce cleaner and more realistic images, meanwhile demanding excessive memory consumption (Lee et al., 2023). The recent works found those representations struggle with low-frequency detail and overfit to high-frequency signals, especially when applying for dynamic scenes, despite using multi-scale representations (Fridovich-Keil et al., 2023; Cao & Johnson, 2023; Peng et al., 2023). While those have marginally had success in the reconstruction of NeRF with the assistance of denoising penalties like total variation (Sun et al., 2023; Fridovich-Keil et al., 2023), they still lack adequate representation of low-frequency spectral features like object shapes and dynamic motion, as shown in 1(a).
To alleviate this issue, we introduce a simple yet powerful approach to fundamentally improve the performance of static and dynamic NeRFs from sparse inputs. In this framework, coordinate-based MLP features handle low-frequency context, while multiple-plane features capture fine-grained details aligning with the spectral bias mentioned earlier. This approach yields three main benefits. First, aligning with the distinct spectral biases of heterogeneous features results in a model less sensitive to hyperparameters and performance variations related to the scene. This is achieved by avoiding both underfitting and overfitting. Second, it allows for stable training through gradual changes in their spectral biases, as discussed in the works by Lin et al. (2021); Yang et al. (2023). Lastly, it facilitates efficient parameter allocation by replacing the need for a low-resolution grid with coordinate-based features.
@@
  • When we reduce the number of parameters, the proposed method still shows competitive performance. This results from skipping the allocation of a spatial low-resolution grid and replacing it with coordinate-based features.
@@ 2 Related Work
Early efforts incorporated pre-trained networks trained on large datasets to compensate for the lack of training data (Jain et al., 2021; Yu et al., 2021; Wang et al., 2021). Another alternative approach incorporated additional information, such as depth or color constraints, to ensure the preservation of 3D coherence (Deng et al., 2022; Yuan et al., 2022; Roessle et al., 2022; Truong et al., 2023). Without the assistance of off-the-shelf models and additional, this line of works devised new regularization to train NeRFs with fewer than ten views. Reg-NeRF incorporates patch-wise geometry and appearance regularization (Niemeyer et al., 2022). This paper verified that their regularization performs well on forward-facing examples like the LLFF dataset. They did not validate object-facing scenes because this assumption demands a high correlation between adjacent views. Recently, progressively manipulating the spectrum of positioning encoding from low to high frequency proved effective in mitigating over-fitting without relying on additional information (Yang et al., 2023; Song et al., 2023). Compared to explicit representations, those still suffer from unsatisfactory visual quality, characterized by blurry boundaries. Recent studies using total variation regularization on explicit representations get rid of artifacts and construct smoother surfaces (Cao & Johnson, 2023; Fridovich-Keil et al., 2023; Sun et al., 2023). However, our findings indicate that this regularization can introduce artificial details that seem real but are not in the data. This can also result in the model failing to converge in certain scenes. We present this problem in the experiments, both qualitatively and quantitatively.
@@
A few works attempted to use explicit parameterization with the sinusoidal encoding of coordinates, but their direction differs from our method since they mainly focus on enriching available features or fewer parameterization, as well as they did not demonstrate the role of tri-planes and coordinate features (Wang et al., 2023b; Peng et al., 2023; Lee et al., 2023). In this paper, our new approach proposes incorporating two distinct features: coordinate-based and multiple-plane features. We emphasize that the disentanglement of these two heterogeneous features is crucial for reliably constructing NeRFs in sparse inputs. The proposed method performs well even with higher-dimensional cases like dynamic NeRFs and extremely limited sparse inputs.
Figure 2: Conceptual illustration of the proposed method utilizing global contexts by coordinate networks and fine-grained details by multi-plane encoding. This method effectively displays two heterogeneous features. The number 1 indicates the use of coordinate network alone, while the symbol 1$+$2 means the use of both coordinated-based MLP network and multi-plane representation.
@@ 3.1 Architecture
where, $\{W_{l},b_{l}\}_{l=1}^{L}$ are the weights and biases of the $l$-layer MLP. In the residual connection blocks, when $l\leq 2$, the block includes two pairs of weights and biases. For $l>2$, the subsequent processes contain one pair of weights and biases. $\oplus$ indicates the concatenation of features.
The output layers use different activations, such as the softplus function for density and the sigmoid function for color. The proposed residual connection allows the network to robustly maintain low-frequency spectral bias from coordinate networks without interference from multi-plane features. Our empirical findings demonstrate that this operation promotes the disentanglement of two features, aligning with a spanning diverse spectrum. A detailed analysis of this residual connection is provided in subsection 4.5 along with the performance gain by this architecture.
@@ 4 Experiments
To answer those questions, we conduct vast experiments over scenarios of two sparse input cases: a few-shot static case and a 4-dimensional dynamic case. To provide a clearer insight into the role of low-frequency representation, we include 2D image regression as an illustrative example. We also include ablation studies to substantiate the rationale behind the architectural choices. We choose the datasets as in-ward-facing object poses, as they are more likely to be occluded by the objects from various viewing locations than forward-facing poses. For performance evaluation, we employ the PSNR metric to gauge image reconstruction quality. In addition, SSIM and LPIPS scores are reported to assess the perceptual quality of the rendered images. Further experimental details are described in Appendix D.
@@ 4.1 Motivation Example: Image Regression
We start by demonstrating the diversity in spectrum information the proposed method possesses in $\{512\times 512\}$ Image Regression tasks. We compare our method with FFN (Tancik et al., 2020) as a sinusoidal embedding and iNGP (Müller et al., 2022), K-Planes (Fridovich-Keil et al., 2023) as explicit parameterization. We also include CAM (Lee et al., 2023), a similar approach that combines coordinate-based networks and grid-based representation for efficient parameterization. Figure 4 shows how baselines handle low and high freqeuncy or resolution features by presenting the average magnitude spectrum by Fourier feature transform. Specifically, in the case of low-frequency features or grids, FFN and CAM, which manipulate their spectrum via sinusoidal encoding, use only the lowest 26 frequencies (10%) of total frequency range. K-Planes, modifed to use four multi-scale tensorial planes, employs only use the lowest resolutional plane with a scale of $32\times 32$. iNGP, with 16 levels of spatial hash-grid, uses only the first two lowest resolutional grids. The proposed method only utilizes the coordinate-based MLP networks with four frequeices $\{2^{i}|i=0,1,2,3\}$. On the other hand, as the finest feature, K-Planes, CAM and Ours all adopt a $128\times 128$ plane.
In Figure 4, we observe that all baselines differs the average magnitude spectrum of rendered images between low and full feature engagement, though the extent of manignitude varies. Low feature engagement describes only low-frequency details, resulting in a minimum spectrum magnitude, whereas full feature engagement captures intrigate details, achieving the higest spectrum magnitude. Examining each instance, the sinusoidal method (FFN) faces an underfitting issue as it overly focuses on low frequency details. Explicit representations like iNGP and K-Planes, however, tend to interpret low resolution features focusing on high frequency details despite only low-resolution features are activated. Surprisingly, CAM, despite incoporating various spectral sinusodial embeddings, also struggles to capture low frequency details. This implies explicit representation such as grid or plane cannot effectively handle low frequency details without careful designs.
@@ 4.3 Dynamic Radiance Fields
4.4 Stability in Sparse-Input NeRFs
In sparse-input NeRFs, stability is defined as the ability to counteract overfitting. We measure stability by evaluating the minimal performance discrepancy between test viewpoints adjacent to and not adjacent to the training views. Specifically, we examine the variance of PSNR across all test viewpoints in the static NeRF dataset. The total variance of PSNR across all images is calculated using 8,000 images from 8 scenes, each with 200 test viewpoints and five trials.
Table 3: Variance of PSNR($\downarrow$) on the static NeRF datasets.
FreeNeRF iNGP TensoRF K-Planes Ours
17.31 23.95 23.22 19.61 18.23
FreeNeRF, which uses MLP and sinusoidal encoding, shows the lowest variance among baselines. Spatially explicit methods like iNGP and TensoRF exhibit significant variances due to their tendency to overfit the training views. While K-Planes reduces instability compared to these methods, its variances still do not match ours. Quantitatively, our method achieves comparable results to FreeNeRF. However, as shown in Table 3, FreeNeRF generally lacks reconstruction performance. Additionally, K-Planes struggles with reconstructing specific scenes, such as ship. In contrast, our method consistently reconstructs all scenes with high quality, avoiding significant degradation. This is also evident in Figure 5. While FreeNeRF exhibits blurry details and K-Planes displays strange color distortion, our method shows the cleanest results without noticeable distortion or artifacts. Considering that our method shows low variances and achieves the highest PSNR, we emphasize the distinction of our approach in terms of both stability and superior capability.
@@ 4.5 Ablation Study
Table 7: Quantitative results between activation and absence of curriculum weighting for multi-plane encoding. Mean and variance of PSNR are presented for each scene. We conduct three trials using random seeds to measure average PSNR and its variance
Metric Static NeRF D-NeRF
Drums Lego Mic
Hell
Warrior
Lego
Stand
up
Trex
CL Mean($\uparrow$) 20.20 26.84 26.61 19.46 24.05 26.22 26.70
Variance($\downarrow$) 2.31 7.20 6.28 2.05 2.93 5.42 3.22
Non- CL Mean($\uparrow$) 19.84 26.54 26.49 19.83 23.94 25.59 26.63
Variance($\downarrow$) 2.55 6.26 9.16 2.80 4.15 4.99 3.40
To validate the curriculum weighting, we conduct a comparison between the proposed method and the same architecture that does not utilize progressive training. We choose {Lego, Drums, Mic} from the static NeRF, and {Hellwarrior, Lego, Standup, Trex} from the dynamic NeRF where this weighting is applied (See Appendix C). We denote CL as the activation of progressive training and Non_CL as its absence. In static NeRFs, we observe that CL consistently has a positive impact on performance improvement to Average PSNR, despite the fact that their improvement on reconstruction is minor, ranging from 0.2 to 0.4 in all cases. For variances, {drums, lego} show no significant difference, but the Mic result indicates that CL mitigates instability, with variances decreasing by 2.8. This reduction means significantly less discrepancy between images adjacent and not adjacent to training views. The effectiveness of progressive training is more pronounced in dynamic NeRFs. While it does not provide significant improvement in the hellwarrior case, it evidently enhances performance in the standup case, leading to 0.6 increase in average PSNR. In terms of stability, we observe that most cases are less sensitive to overfitting with progressive training. Although variance slightly increases in the standup, it is not significant issue considering average PSNR improvement. In summary, progressive training influences on either performance improvement or mitigation of instability in sparse-input NeRF by gradually engaging multi-plane channels. This allows the initial channels to learn global details, while later channels focus more on finer details. The detailed explanation, including graphics, is provided in Appendix H.
@@ 5 Conclusion
In this paper, we introduce refined tensorial radiance fields that seamlessly incorporate coordinate networks. The coordinate network enables the capture of global context, such as object shapes in the static NeRF and dynamic motions in the dynamic NeRF dataset. This property allows multi-plane encoding to focus on describing the finest details. Through extensive experiments, we demonstrate that the proposed method consistently outperforms the baselines and their regularization in the few-shot regime. Notably, the proposed method exhibits strong stability, showing less discrepancy between images adjacent and non-adjancet to training views. Additionally, it preserves performance even with a reduced number of parameters.
Impact Statement
@@ C.1 Hyper-parameters on the Static NeRF
The proposed model incorporates multi-plane encoding and MLPs with skip connections. For the multi-plane encoding, we utilize 48-dimensional channels. The resolution of plane features is upscaled to 8,000,000 ($200^{3}$) by the end of training. The hyper-parameters, such as the weight for Laplacian smoothing ($\lambda_{1}$), the curriculum learning schedule, and the initial feature resolution, vary across scenes as part of our hyperparameter tuning to achieve optimal results. However, we use same $\lambda_{2,3}$ across scenes. In this dataset, $\lambda_{2}=1$ is used, and $\lambda_{3}$ is initially set to $0.00008$ at the start of training, increasing to $0.00004$. This is the same weighting strategy used in the previous method (Chen et al., 2022). Detailed information on the hyperparameters for multi-plane encoding can be found in Table C.1. For the decoder, we employ standard fully connected layers with ReLU activations, each containing 256 channels. The encoder consists of four fully connected ReLU layers, with a skip connection introduced after the second layer, which concatenates the fused input features. Occupancy is directly calculated from the obtained features with softplus function applied to the first channel. The RGB decoder, following this, consists of two layers. The color values are obtained from the features processed by the RGB decoder through sigmoid activation.
@@ C.2 Hyper-parameters on the Dynamic NeRF
The configuration for the dynamic Neural Radiance Fields (NeRF) case adheres to the same settings as the static case. We utilize plane features with 48 channels. The initial voxel resolution is set at 4,096 ($16^{3}$) and is subsequently upscaled to 8,000,000 ($200^{3}$). For hyper-parameters, $\lambda_{2}$ is set to 2.5 for all scenes in the dynamic NeRFs to ensure smoother multi-plane features along the time-axis. This approach is proven effective in previous work (Cao & Johnson, 2023). The parameter $\lambda_{3}$ is set to 0.00001 across all dynamic NeRF scenes. For a more detailed description, please refer to Table C.2. The structure of the decoder, initial learning rate, and optimizer configuration remain identical to those used in the static NeRF. Any configurations not specified here follow directly from the HexPlane method as described in (Cao & Johnson, 2023).
@@ Appendix E The Evaluation Statistics of Static and Dynamic NeRF Datasets
Variance of PSNR on the static NeRF datasets.
We elaborate on the variance of PSNR for each instance in the static NeRF dataset in Table E.7. Specifically, we examine the variance of PSNR across all test viewpoints in the static NeRF dataset. The total variance of PSNR across all images is calculated using 8,000 images from 8 scenes, each with 200 test viewpoints and five trials.
Examining Table E.7, while FreeNeRF records the lowest variance across nearly all scenes except for Ficus and Mic, it consistently lags behind the top score across all scenes. This observation highlights FreeNeRF tends to underfit. Notably, in scenes like Ficus and Mic where caputring intricate and delicate signals such as thin leaves or mesh structures is crucial, FreeNeRF struggles. This suggest that FreeNeRF is stable, but it faces difficulties in effectively rendering complex details effectively. On the other hand, methods like iNGP and TensoRF exhibit poor average PSNR scores such as Drums and Ship, and achieve low variances. This pattern indicates a failure to learn effectively, resulting in meaningless output and negligible difference between training and test views due to overfitting. Although these methods sometimes perform well in scenes with simple geometry like Hotdog and Lego, they often lack stability and underperform in scenarios requiring sparse input handling for NeRF applications.
However, as stated in subsection 4.4, our method consistently shows low variances in most scenes—except for hotdog and materials, where we achieve the highest PSNR—we emphasize the distinction of our approach in terms of both stability and superior capability compared to K-Planes.
Table E.7: Variance of PSNR($\downarrow$) on the static NeRF dataset.
Chair Drums Ficus Hotdog Lego Materials Mic Ship Total
FreeNeRF 5.07 1.72 8.72 11.16 6.42 2.13 17.28 6.48 17.31
iNGP 8.43 3.13 1.37 12.39 7.78 2.96 9.02 6.03 23.95
TensoRF 10.88 2.86 2.17 13.11 10.27 2.82 8.06 5.71 23.22
K-Planes 10.74 2.55 2.83 27.19 10.76 2.99 9.23 11.48 19.61
Ours 3.82 2.82 2.00 16.38 8.72 4.09 9.68 6.01 18.23